OpenStudy (anonymous):

use polar coordinates to find the volume of a sphere of radius a. Help, please

OpenStudy (anonymous):

I set up theta from 0 to 2pi and r from 0 to a , the function is 2pi r^2 but the result is not right.:(

terenzreignz (terenzreignz):

Everything always falls into place once the integral is set up.

OpenStudy (anonymous):

@terenzreignz what do you mean?

Parth (parthkohli):

Terence is epic.

Parth (parthkohli):

Integrate the function.

terenzreignz (terenzreignz):

^But to integrate a function, you need to have... A FUNCTION Now tell me, @Hoa What is the equation of a sphere of radius A, centred around the origin?

terenzreignz (terenzreignz):

Sorry, radius a.

Parth (parthkohli):

:')

OpenStudy (anonymous):

2pir^2

OpenStudy (anonymous):

sorry a^2

terenzreignz (terenzreignz):

Hardly the equation of a sphere, @Hoa But don't worry, that's why I'm here :) The equation of a sphere centred around the origin, with radius a, is \[\LARGE a^2=x^2+y^2+z^2\]

OpenStudy (anonymous):

yeap

terenzreignz (terenzreignz):

You'll notice it's similar to the equation of a circle right? In fact, we might as well put that up... \[\LARGE r^2=x^2+y^2\]

terenzreignz (terenzreignz):

Now, even though this isn't the correct approach, when converting from rectangular coordinates to cylindrical, you use this relation, as written above, right? ^ You need to change coordinate systems from \[\LARGE (x,y,z) \rightarrow (r,\theta,z)\] Fortunately, and clearly, z remains unchanged, so it's only a matter of the first two coordinates.

terenzreignz (terenzreignz):

Conveniently, in the equation of the sphere with radius a, the part \(\large x^2+y^2\) is already present. \[\LARGE a^2=x^2+y^2+z^2\]

terenzreignz (terenzreignz):

But first things first.. you've already set up the limits of the function, but now, we have to find THE function. It's almost always a function z = f(x,y). Can you express z as a function of x and y? (Remember that a is constant). @Hoa

OpenStudy (anonymous):

x^2 +y^2 =a^2

terenzreignz (terenzreignz):

^That is not a function. What I'm asking you to do is to isolate z, and express it as a function of x and y. Can you do that?

OpenStudy (anonymous):

you mean z=sqr(a^2 -x^2 -y^2)?

terenzreignz (terenzreignz):

Much better.

OpenStudy (anonymous):

+-

terenzreignz (terenzreignz):

\[\huge z = f(x,y) = \sqrt{a^2-(x^2+y^2)}\]

OpenStudy (anonymous):

why don't we take - value?

terenzreignz (terenzreignz):

No, we are not taking +/- values. Just the positive values.

OpenStudy (anonymous):

because the value of a volume must be a positive number?

terenzreignz (terenzreignz):

We can do that, but that complicates things, don't you think? If we take only the positive values, we are only taking the volume of the upper hemisphere of the sphere. In other words, only half the sphere. This will be rectified when we get to the integral. We are only taking positive values because we need a FUNCTION, which, as you know, may not have more than one value, for any specific x-value or y-value.

OpenStudy (anonymous):

and then *2?

terenzreignz (terenzreignz):

That is correct. You catch on quick. ;)

terenzreignz (terenzreignz):

So, we have a function. But we need it to be a function in polar coordinates... How on earth do you change a function f(x,y) into a function g(r,θ) ?

OpenStudy (anonymous):

by using x =rcos theta y = r sin theta?

terenzreignz (terenzreignz):

Well, as I said earlier, the expression \(\large x^2 + y^2\) appears, quite conveniently, and may readily be converted to \(\large r^2\)

terenzreignz (terenzreignz):

Well, you could do that, but you'll get the same result... after a little more crunching. ;) Best keep things simple.

OpenStudy (anonymous):

ok, so the function is sqr (a^2 -r^2)?

terenzreignz (terenzreignz):

That is correct.

terenzreignz (terenzreignz):

So, now, we integrate our function \[\huge \iint\limits_R \sqrt{a^2-r^2} \ dA\]

terenzreignz (terenzreignz):

You've defined the region over which we integrate, right?

OpenStudy (anonymous):

yes, sir . everything done after set up the limits as I do above, right?

OpenStudy (anonymous):

hey, hold on, how about one more r?

terenzreignz (terenzreignz):

No calling me sir. ;) Let's set up the limits... \[\huge \int\limits_{\theta=0}^{2\pi}\int\limits_{r=0}^{a}\sqrt{a^2-r^2} \ dA\]

hartnn (hartnn):

just a note, \(\huge V=2\iint\limits_R \sqrt{a^2-r^2} \ dA \\ dA =r dr d \theta\)

terenzreignz (terenzreignz):

Now, in polar coordinates, you can just take this to faith, unless you intend to take Real Analysis... but \[\huge dA=dxdy=r\cdot drd\theta\]

OpenStudy (anonymous):

@hartnn got it.

OpenStudy (anonymous):

Thanks terenzeignz, thanks a lot everybody, I am ok from then.

terenzreignz (terenzreignz):

So, the integral becomes \[\huge \int\limits_{\theta=0}^{2\pi}\int\limits_{r=0}^{a}\ r \sqrt{a^2-r^2} \ drd\theta\ \]

OpenStudy (anonymous):

bingo.

terenzreignz (terenzreignz):

You can do it from here?

OpenStudy (anonymous):

sure.

terenzreignz (terenzreignz):

If you say so. Good Hunting :)

OpenStudy (anonymous):

do you want to check?

OpenStudy (anonymous):

let u = inside sqr, change the limit. and... done, right?

terenzreignz (terenzreignz):

I'll leave you to your methods. :)

OpenStudy (anonymous):

I do shortcut, but trust me, I can do it. thanks a lot.

terenzreignz (terenzreignz):

Though, you're welcome to show your process.

OpenStudy (anonymous):

under those pressure?

terenzreignz (terenzreignz):

Pressure is an illusion... They are not there to intimidate you, they are there... to cheer you on. Get inspired. And did I mention... don't call me sir? Good :)

OpenStudy (anonymous):

ok, thanks friend!!!

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

to me, u = a^2 -r^2 ---> du = -2 r dr , substitute into the function to get -1/2sqr (u) du

OpenStudy (anonymous):

change the limit , and take int as usual, it's quite tedious but it is easy

terenzreignz (terenzreignz):

Stand by while I attempt to locate a sheet of paper and a pencil.

OpenStudy (anonymous):

at the end up, the volume of a sphere must be 4/3 pi r^3

terenzreignz (terenzreignz):

a, not r.

OpenStudy (anonymous):

yes. sorry. a^3

OpenStudy (anonymous):

ha ha.. you got 5 medals from my question, cheeerr !!!

terenzreignz (terenzreignz):

Isn't that just great :) Here, take mine ...

OpenStudy (anonymous):

haha...ridiculous us!!! cheer up each other!!! ok anyways, I appreciate what you do for me. take care

terenzreignz (terenzreignz):

Could you restate your new question?

OpenStudy (anonymous):

find the volume of the solid above z= sqr(x^2 +y^2) and below the sphere x^2 +y^2+z^2 =1. I confused

OpenStudy (anonymous):

hehe... I am safe now, no one monitor, no pressure.

terenzreignz (terenzreignz):

You could do this normally, with rectangular coordinates, but I think polar is the way to go here, considering the ease of integration. Do you agree with me on that count? :)

OpenStudy (anonymous):

yeap

terenzreignz (terenzreignz):

Well, remember in 2d, when you have two functions and are asked to find the area between them  You simply integrate the function f(x) - g(x) right?